OpenStudy (anonymous):

Write the general form of the equation of each circle. Center at the point (4,6) and tangent to the x-axis.

OpenStudy (anonymous):

ok then i go to bed

OpenStudy (anonymous):

Lol. Okay. :]

OpenStudy (anonymous):

ok that was wrong \[(x-4)^2+(y-6)^2=r^2\]

OpenStudy (anonymous):

now where do you think you will find \(r\) ?

OpenStudy (anonymous):

i.e. how far is \((4,6)\) from the \(x\) axis? (don't think too hard)

OpenStudy (anonymous):

want a hint?

OpenStudy (anonymous):

6 away?

OpenStudy (anonymous):

Yes please!

OpenStudy (anonymous):

you don't need one clearly

OpenStudy (anonymous):



OpenStudy (anonymous):

so yeah it is 6 units away final answer \[(x-4)^2+(y-6)^2=36\]

OpenStudy (anonymous):

oh! Thank you! <33

OpenStudy (anonymous):

yw and now @Wondergiirl you are on your own g'night

OpenStudy (anonymous):

Night ! :D